OpenStudy (anonymous):

Factor. 4x^4-19x^3-5x^2

OpenStudy (anonymous):

\[4x ^{4}-19x ^{3}-5x ^{2}\]

OpenStudy (anonymous):

\[4x^4-19x^3-5x^2=x^2(4x^2-19x-5)\] is a good start

OpenStudy (anonymous):

Then factor by grouping?

OpenStudy (anonymous):

i don't know i suck at factoring as far as i can tell factoring by grouping is the same as factoring by trial and error

OpenStudy (anonymous):

Haha :) Well thank you very much, I think I've got it by now. I suck at factoring when it comes to anything to a degree higher than 2

OpenStudy (anonymous):

i would set it equal to zero and solve the answer will be \[x^2(4x+1)(x-5)\]

OpenStudy (anonymous):

how you factor \[4x^2-19x-5\] using any method other than trial and error is a mystery to me if you have a method, bless you

OpenStudy (anonymous):

or quadratic formula but for this you will need to get the equation in (ax^2+bx+c=0) form then its easier and it would give you answers x = 0, x= -1/4 and x= 5

OpenStudy (anonymous):

Well, from what you already did for me (sorry for saying it like that but you did) All I gotta do now is multiply A and C and find factors of B that multiply to the product of AC and add up to B

OpenStudy (anonymous):

@satellite73 i have a question

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@lankymonkey i don't answer questions

OpenStudy (anonymous):

how did you get to smartscore 100

OpenStudy (anonymous):

oh umm sorry satelite are you a robot?

OpenStudy (anonymous):

?????????????? Private message each other ????????????

OpenStudy (anonymous):

umm ok sry satelite

OpenStudy (anonymous):

why was satelite acting weird? and why do moderators and admins hate me????